Professor Counts his midterms as 2/3 of the grade and his final as 1/3of the grade. If Wendy scored 48 on the midterm, What range of scores on the final exam would put Wendy’s average between 70-79 inclusive?
Not from a book or ISBN#

Answer by nerdybill(7384)

Let x = final grade
.
The inequality looks like this:
70 <= (2/3)(48) + (1/3)x <= 79
70 <= (2)(16) + x/3 <= 79
70 <= 32 + x/3 <= 79
.
Looking at the left inequality first:
70 <= 32 + x/3
210 <= 96 + x
114 <= x
.
Looking at the right inequality:
32 + x/3 <= 79
96 + x <= 237
x <= 141
.
Putting it all together:
114 <= x <= 141
This means Wendy has to score between 114 and 141 on her final exam.